40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 13017     Entry time: Mon May 29 16:47:38 2017     In reply to: 13015     Reply to this: 13019
Author: gautam 
Type: Update 
Category: General 
Subject: Coil driver boards reinstalled 

Yesterday, I reinstalled the de-whitening boards + coil driver boards into their respective Eurocrate slots, and reconnected the cabling. I then roughly re-aligned the ITMs using the green beams. 

I've given Steve a list of the thin-film resistors we need to implement the changes discussed in the preceeding elogs - but I figured it would be good to see if we can realize the projected improvement in MICH displacement noise just by fixing the BS Oplev loop shape and turning the existing whitening on. Before re-installing them however, I did make a few changes:

  • Removed the gain of x3 on all the signal paths on the De-Whitening boards, and made them gain x1. For the De-Whitened path, this was done by changing the feedback resistor in the final op-amp (OP27) from 7.5kohm to 2.49kOhm, while for the bypass path, the feedback resistor in the LT1125 stages were changed from 3.01kohm to 1kohm. 
  • To recap - this gain of x3 was originally implemented because the DACs were +/- 5V, while the coil driver electronics had supply voltage of +/- 15V. Now, our DACs are +/- 10V, and even though the supply voltage to the coil driver boards is +/- 15V, in reality, the op-amps saturate at around 12V, so we aren't really losing much in terms of range.
  • I also modified the de-whitening path in the BS de-whitening board to mimic the configuration on the ITM de-whitening boards. Mainly, this involved replacing the final stage AD797 with an OP27, and also implementing the passive pole-zero network at the output of the de-whitened path. I couldn't find capacitors similar to those used on the ITM de-whitening boards, so I used WIMA capacitors.
  • The SRM de-whitening path was not touched for now.
  • On all the boards, I replaced any AD797s that were being used with OP27s, and simply removed AD797s that were in DAQ paths.
  • I removed all the potentiometers on all the boards (FAST analog path on the coil driver boards, and some offset trim Pots on the BS and SRM de-whitening boards for the AD797s, which were also removed).
  • For one signal path on the coil driver board (ITMX ch1), I replaced all of the resistors with thin-film ones and re-measured the noise. However, the excess noise in the measurement below ~40Hz (relative to the model) remained.

Photos of all the boards were taken prior to re-installation, and have been uploaded to the 40m Google Photos page - I will update schematics + photos on the DCC page once other planned changes are implemented.

I also measured the transfer functions on the de-whitened signal paths on all the boards before re-installing them. I then fit everything using LISO, and updated the filter banks in Foton to match these measurements - the original filters were copied over from FM9 and FM10 to FM7 and FM8. The new filters are appended with the suffix "_0517", and live in FM9 and FM10 of the coil output filter banks. The measured TFs (for ITMs and BS) are summarized in Attachment #1, while Attachment #2 contains the data and LISO file used to do the fits (path to the .bod files in the .fil file will have to be changed appropriately). I used 2 complex pole pairs at ~10 Hz, two complex zero pairs at ~100Hz, real poles at ~15Hz and ~3kHz, and real zeros at ~100Hz and ~550Hz for the fits. The fits line up well with the measured data, and are close enough to the "expected" values (as calculated from component values) to be explained by tolerances on the installed components - I omit the plots here. 

After re-installing the boards in the Eurocrate, restoring rough alignment, and updating the filter banks with the most recent measured values, I wanted to see if I could turn the whitening on for one of the optics (ITMY) smoothly before trying to do so in the full DRMI - switching off the "SimDW_0517" filter (FM9) should switch the signal path on the de-whitening board from bypass to de-whitened, and I had confirmed last week with an extender board that the voltage at the appropriate backplane connector pin does change as expected when the FM9 MEDM button is toggled (for both ITMs, BS and SRM). But today I was not able to engage this transition smoothly, the optic seems to be getting kicked around when I engage the whitening. I will need to investigate this further. 


Unrelated to this work: the ETMY Oplev HeNe is dead (see Attachment #3). I thought we had just replaced this laser a couple of months ago - what is the expected lifetime of these? Perhaps the power supply at the Y-end is wonky and somehow damaging the HeNe heads?

Attachment 1: deWhitening_consolidated.pdf  148 kB  Uploaded Mon May 29 17:50:00 2017  | Hide | Hide all
deWhitening_consolidated.pdf deWhitening_consolidated.pdf deWhitening_consolidated.pdf
Attachment 2: deWhitening_measurements.zip  99 kB  Uploaded Mon May 29 17:51:18 2017
Attachment 3: ETMY_OL.png  25 kB  Uploaded Mon May 29 17:52:00 2017  | Hide | Hide all
ETMY_OL.png
ELOG V3.1.3-